# PaTAS: A Parallel System for Trust Propagation in Neural Networks Using Subjective Logic

### v2

@@ V-A2 Output-Trust Aggregation
The Trust Feedforward process produces a vector of trust values, one for each output neuron. The goal of Output-Trust Aggregation is to combine these individual opinions into a single, consolidated trust score representing the overall trust in the network’s output. This combination is done by means of a fusion operator. Alternatively, instead of aggregating across all outputs, one may directly consider the trust assigned to the final decision. For instance, in digit classification, if the model outputs a probability $0.9$ for class ‘1’, the trust in the decision can be taken as the trust score associated with the output neuron for class ‘1’.
V-A3 GenIPTA
@@
The purpose of the GenIPTA function is to dynamically construct a function tailored to a specific inference. This function, called Inference-Path Trust Assessment (IPTA), reflects how trustworthy the exact computational path taken during inference is. When an inference is performed, contextual information is recorded, such as the list of neurons activated along the path. In this case, the activation trace is used to instantiate a temporary subnetwork of Trust Nodes containing only the trust nodes corresponding to those activations, thereby mirroring the precise inference path of the neural network.
@@
Contextual information is not limited to activation traces. Typically, all neurons are activated to some degree, but only a subset of these activations is strongly relevant to the decision. GenIPTA can therefore operate on truncated activation sets that retain only the most relevant neurons, which allows for more focused and interpretable trust assessments of the actual decision-making process. GenIPTA can also be more complex and take into account the specific activation values while keeping all neurons, performing a weighted trust assessment where stronger activations contribute more heavily to the inference.
V-B Parameter-Trust Revision
@@
2:  Summary: Revises and updates the trust parameters of the Trust Nodes Network by combining gradient evidence, label trust, and neuron input trust, under mini-batch training.
@@
7:     Step 2: Gather gradient evidence for neuron $n_{i}^{(l)}$
@@
The design of Parameter-Trust Update is motivated by the standard backpropagation algorithm used in neural network training. In backpropagation, the gradient of the loss function with respect to the network parameters indicates how much each parameter contributes to the output error. Thus, for a parameter $\theta$, the update is given by
@@
The Parameter-Trust Update process is formally described in Alg. 1. For each layer in the network, the system revises the trust parameters of every trust node based on the evidence available from the training batch, the input processed during feedforward, and the gradients. It then updates these trust values to reflect the reliability of the newly updated parameters, using evidence from the learning rate and the intermediate trust scores.
The algorithm runs once per batch and begins by computing a combined trust opinion over all labels in the current batch:
@@
For each neuron $n_{i}^{(l)}$ at index $i$ in layer $l$, the system computes:
  • $T_{n_{i}^{(l)}\mid y_{\text{batch}}}$, the trust conditioned on the current batch labels. This is inferred by checking each incoming weight gradient $g^{(l)}_{i,j}$: if $|g^{(l)}_{i,j}|<\epsilon$ it is counted as positive evidence $r$, otherwise as negative evidence $s$. These counts are then mapped into a binomial opinion using the Baseline-Prior Quantification model.
@@
    Revision with deduced trust using fusion:
@@
    Adjustment with auxiliary factors: During training, the update of parameter $\theta_{i,j}^{(l)}$ depends not only on its current value but also on auxiliary factors such as the learning rate $l_{r}$, the input feature $x^{(l-1)}_{j}$, and the label $y_{\text{batch}}$ (see Eqs. 15 and 16). To reflect this, we adjust parameter trust using:
@@
    This formulation reflects the fact that both unreliable features and mislabeled data can strongly bias parameter updates. Although it might introduce some redundancy, it ensures that trust in both input features and labels is explicitly propagated into the parameter trust.
@@
V-C Theoretical Properties of PaTAS
V-C1 Convergence of PaTAS
The convergence of the PaTAS is governed by the stability of its inputs and the structure of its recursive Parameter-Trust Update process. During training, PaTAS updates the internal trust opinions associated with network parameters using trust signals from the inputs, labels, hyperparameters, and gradient information. This update mechanism is outlined in Alg. 1. As proved in Theorem 3, the PaTAS converges under specific situation. This convergence lies on some specific characteristic of the operators used to feedforward and revise the trust in the parameters $\theta$.
Before delving into the specifics of PaTAS convergence, we first present a key theorem that will serve as a foundational tool in establishing the convergence properties of PaTAS.
Theorem 1 (Convergence of Subjective Logic Arithmetic Sequence).
Let $(\Omega,\ominus)$ a group. $\Omega$ is set of opinions which can be seen as $0,1]^{4}$ Let $\omega_{n}=(b_{n},d_{n},u_{n},a_{n})\in\Omega$ be a sequence defined recursively by the operator $\ominus$ as follows:
$\omega_{n+1}=\omega_{n}\ominus q,$
where $q\in\Omega$ and the operator $\ominus$ is a fusion operator.
Then the sequence $(\omega_{n})$ converges in $0,1]^{4}$ to a limit:
Proof.
See Appendix C
Theorem 2 (Convergence of Subjective Logic Geometric Sequence).
Let $(\Omega,\odot)$ a group. $\Omega$ is set of opinions which can be seen as $0,1]^{4}$ Let $\omega_{n}=(b_{n},d_{n},u_{n},a_{n})\in\Omega$ be a sequence defined recursively by the operator $\odot$ as follows:
$\omega_{n+1}=\omega_{n}\odot q,$
  • $a_{n}=a_{0}a_{q}^{n}\to 0$ as $n\to\infty$,
  • $d_{n}\to 1$ if $d_{q}>0$, and $d_{n}=d_{0}$ if $d_{q}=0$,
  • $b_{n}\to 0$ if $p_{q}=b_{q}+a_{q}u_{q}<1$, and $b_{n}=b_{0}$ if $p_{q}=1$.
Proof.
See Appendix C
@@ V-C1 Convergence of PaTAS
Let $T_{\theta}^{(n)}$ denote the trust opinion at iteration $n$ for parameter $\theta$. If the revision of the trust in the weights is performed as specified in Alg. 1, then the PaTAS parameter will also converges.
@@
Moreover, the convergence of $T_{x}$, $T_{y}$, and $T_{l_{r}}$ implies that trust inputs to the update mechanism are stable. Consequently, each new trust update for $T_{\theta}$ is computed using consistent and bounded evidence, which over time results in the stabilization of the subjective opinions assigned to $T_{\theta}$.
In detail:
  • $g$ converges so $T_{\theta\mid y_{\text{batch}}}$ will remain the same.
  • since $T_{\theta\mid\mid{Y_{\text{batch}}}}$ is a deterministic calculation from converges $T_{\theta\mid y_{\text{batch}}}$ and $T_{y_{batch}}$ therefore $T_{\theta\mid\mid{Y_{\text{batch}}}}$ also converges
[…]
  • The function $f_{upd}$ in our implementation is based $T_{\theta}$, $T_{l_{r}}$ and $T_{x}$} which all converge.
V-C2 Symmetry and Invariance Properties of PaTAS
@@
Let $x=(b,d,u)$ be a binomial opinion with belief $b$, disbelief $d$, and uncertainty $u$, where $b+d+u=1$. The opinion $\bar{x}=(d,b,u)$ is called the symmetric of $x$. Two binomial opinions $x$ and $\bar{x}$ are symmetric if they share the same uncertainty and have inverted belief and disbelief, i.e.,
$x=(b,d,u),\quad\bar{x}=(d,b,u)\quad\text{with }b+d+u=1.$
@@ VI Evaluation and Results
The goal of our evaluation is to validate the Parallel Trust Assessment System (PaTAS) both theoretically and empirically. Specifically, we aim to demonstrate that PaTAS produces interpretable trust estimates that (i) converge during training under specific conditions, (ii) respect symmetry and invariance properties, and (iii) are able to provide interpretable assessments under realistic conditions such as noisy features, corrupted labels, or adversarial perturbations.
Our evaluation approach combines controlled synthetic degradations with real-world datasets. We systematically vary the trustworthiness of inputs ranging from fully trusted, fully uncertain to fully distrusted and observe how created PaTAS propagates these trust assessments through the network. For each scenario, we track three complementary metrics: trust mass, uncertainty mass, and distrust mass. We also compare these values against standard model accuracy to understand how input trust affects output reliability.
We conduct three experiments of increasing complexity:
@@ VI-A1 Experiment 1 - Breast Cancer Classification
Training Data trust is assessed with three trust profiles for input features and label: fully trusted $(1,0,0)$, fully distrusted $(0,1,0)$, and fully uncertain $(0,0,1)$. These are combined (for inputs and label assessment) to form nine combinations. Data features and labels trust degradation may arise in practice from poor-quality imaging, human annotation errors, or flaws in the preprocessing pipeline. In all experiments, degradations are simulated using controlled perturbation functions. For fully uncertain opinions, we introduce additive uniform noise
$x^{\prime}=x+\epsilon,\quad\epsilon\sim U(-a,a),$
@@
$a=0.3\times\max(\text{feature}).$
Each feature is perturbed with probability $0.3$. After noise addition, if $x^{\prime}$ lies outside the valid range $\min(\text{feature}),\max(\text{feature})]$, it is clipped to the corresponding boundary:
$x^{\prime}=\begin{cases}\min(\text{feature}),&x^{\prime}<\min(\text{feature}),\\ \max(\text{feature}),&x^{\prime}>\max(\text{feature}),\\ x^{\prime},&\text{otherwise}.\end{cases}$
To model distrust, we instead corrupt input features entirely.
Similarly, label degradation is simulated by randomly replacing labels (or replacing all labels for distrust). In addition to these extreme cases, we also consider two intermediate scenarios:
@@
    the same scenario for fully uncertain input features and fully uncertain labels, but where the assessment is set to $(0.25,0.25,0.5)$, reflecting partial distrust and trust, instead of fully uncertain opinion (0,0,1).
@@ VI-A2 Experiment 2 - MNIST
Although these architectures are not state-of-the-art for MNIST, they are sufficient to evaluate the behavior of PaTAS across different model sizes and to demonstrate how trust propagates through the network. Each model uses the ReLU activation function for the hidden layer and Softmax for the output. Models are trained for 10 epochs with a batch size of 18 and a learning rate of 0.001, achieving test accuracies of 89%. We evaluate using fully uncertain Training data trust assessment functions for both input features and labels.
@@ VI-A3 Experiment 3 - Poisoned MNIST
Implementation Details
During inference, multiplication operations in the feedforward phase are implemented using SL trust discounting, while addition operations employ a generalized SL averaging fusion operator to support summations over multiple inputs. Trust revision uses the same averaging fusion. The PaTAS framework and primary neural network were implemented in Python 3.9 using NumPy. Two main modules were developed: PrimaryNN, which handles network structure, training, and inference, and PaTAS, which implements trust assessment and propagation functions defined in the operational flow (Fig. 6). These modules interact to dynamically evaluate and update trust during feedforward and backpropagation. The full implementation, including all modules, dependencies, and experiment scripts, is available as supplementary materials, with detailed instructions for reproducing all experiments.
@@ VI-B Results and Analysis
All results are depicted in Figs. 11, 12, 9, 16, 18, 19, 13, 17, 14 and 15 (Appendix D) and summarized in Tables II, II, III and IV. They confirm several theoretical properties of PaTAS proven in Section V-C:
Trust Assessment $\epsilon=0.01$ $\epsilon=0.1$ Train (%) Test (%)
$X$ fully distrusted, $Y$ distrusted 0 0 63 38
$X$ fully distrusted, $Y$ fully uncertain 0 0 53 62
$X$ fully distrusted, $Y$ fully trusted 0 0 63 63
$X$ fully uncertain, $Y$ distrusted 0 0 96 3
$X$ fully uncertain, $Y$ fully uncertain 0.28 0.31 53 46
$X$ fully uncertain, $Y$ fully trusted 0.30 0.34 97 96.49
$X$ fully trusted, $Y$ distrusted 0 0 98 0.88
$X$ fully trusted, $Y$ fully uncertain 0.27 0.33 52 89
$X$ fully trusted, $Y$ fully trusted 0.79 0.90 98 99
(0.25, 0, 0.75) - 0.38 71 90
(0.25, 0.25, 0.5) - 0.17 53 46
@@
Hidden Neurons Trust mass $t$ Train (%) Test (%)
@@
For detailed results, we primarily focus on the three plots corresponding to fully trusted input. Processing uncertain inputs naturally yields uncertain outputs, while symmetry implies that trusted and distrusted cases mirror each other. We also note that when processing fully trusted inputs, the distrust mass generally remains close to zero. Since trust, uncertainty, and distrust sum to 1, analyzing only the trust mass is sufficient in such cases.
@@
We evaluated the system for two values of $\epsilon$ (0.1 and 0.01). we recall that $\epsilon$ is the threshold used by the NodeTrust function in Alg. 1.
@@
As shown in Table II, when features and labels are noised, accuracy improves as model size increases. However, training accuracy remains consistently lower than test accuracy, with the gap widening for larger networks. This indicates that the models learn useful representations but cannot fully fit noisy training data, while still generalizing well on clean test samples. Test accuracy greater than training accuracy across all architectures, suggesting training noise makes the model underestimate its learning progress. Larger architectures yield more stable trust assessments, likely due to smaller average gradient magnitudes.
@@
In this experiment, PaTAS successfully distinguishes clean from poisoned labels even when corruption is moderate. Trust mass grows rapidly for clean and poisoned labels. However, poisoned labels remain relatively low, demonstrating effective detection of untrustworthy labels (Table III). Patch size strongly affects this separation:
@@
Patch size Trust mass for 3 Trust mass for 6 Train (%) Test (%) Clean 3 Clean 6 Poisoned 3 Poisoned 6

### v3

@@ V-B2 Output-Trust Aggregation
The Trust Feedforward function produces a vector of trust values, one for each output neuron. The goal of Output-Trust Aggregation is to combine these individual opinions into a single, consolidated trust score representing the overall trust in the network’s output. This combination is performed using a SL fusion operator, which fuses the trust opinions of all output Trust Nodes into one aggregated opinion. Alternatively, instead of aggregating across all outputs, one may directly consider the trust assigned to the final decision. For instance, in digit classification, if the model outputs a probability $0.9$ for class ‘1’, the trust in the decision can be taken as the trust score associated with the output neuron for class ‘1’.
V-B3 Inference-Path Trust Assessment Generation(GenIPTA)
@@
The purpose of the GenIPTA module is to dynamically construct a function tailored to a single, specific inference. This function, called Inference-Path Trust Assessment (IPTA), reflects how trustworthy the exact computational path taken during that inference is. When an inference is performed, contextual information is recorded, such as the list of neurons activated along the path. In this case, the activation trace is used to instantiate a temporary subnetwork of Trust Nodes containing only the Trust Nodes corresponding to those activations, thereby mirroring the precise inference path of the neural network.
@@
Contextual information is not limited to activation traces. Typically, all neurons are activated to some degree, but only a subset of these activations is strongly relevant to the decision. GenIPTA can therefore operate on truncated activation sets that retain only the most relevant neurons, which allows for more focused and interpretable trust assessments of the actual decision-making process. The GenIPTA can also work in a more complex way by using the actual activation values of all neurons, performing a weighted trust assessment where stronger activations contribute more heavily to the inference.
The detailed functional flow of PaTAS across feedforward, backpropagation, and inference is illustrated in Fig. 7.
V-C Parameter-Trust Update
@@
2:  Summary: Revises and updates the trust parameters of the Trust Nodes by combining gradient evidence, label trust, and neuron input trust, under mini-batch training.
@@
7:     Step 2: Gather gradient evidence $g$ for neuron $n_{i}^{(l)}$
@@
The notation used throughout this subsection and in the Algorithm 1 is summarized in Table VII in Appendix A.
The design of the Parameter-Trust Update is motivated by the standard backpropagation algorithm used in neural network training. In backpropagation, the gradient of the loss function with respect to the network parameters indicates how much each parameter contributes to the output error. Thus, for a parameter $\theta$, the update is given by
@@
The Parameter-Trust Update process is formally described in Alg. 1. For each layer in the network, the framework revises the trust parameters of every Trust Node based on the evidence available from the training batch, the input processed during feedforward, and the gradients. It then updates these trust values to reflect the reliability of the newly updated parameters, using evidence from the learning rate and the intermediate trust scores.
The algorithm runs once per batch and begins by computing a combined trust opinion over all labels in the current batch (Line 3):
@@
For each neuron $n_{i}^{(l)}$ at index $i$ in layer $l$, the framework computes:
  • $T_{n_{i}^{(l)}\mid y_{\text{batch}}}$ (Line 9), the trust conditioned on the current batch labels. This is inferred by checking each incoming weight gradient $g^{(l)}_{i,j}$: if $|g^{(l)}_{i,j}|<\epsilon$ [^2] The threshold $\epsilon$ is not a tuned hyperparameter but a sensitivity parameter used only in the NodeTrust function to distinguish weak from strong gradients. Its value can be chosen in different reasonable ways depending on the desired sensitivity; in this work, we select $\epsilon$ relative to the typical gradient scale during training. Importantly, $\epsilon$ does not affect model predictions or training dynamics, but only the sensitivity of the trust-update mechanism. it is counted as positive evidence $r$, otherwise as negative evidence $s$. These counts are then mapped into a binomial opinion using the Baseline-Prior Quantification model.
@@
    Revision with deduced trust using fusion (Line 14):
@@
    Adjustment with auxiliary factors (Line 16): During training, the update of parameter $\theta_{i,j}^{(l)}$ depends not only on its current value but also on auxiliary factors such as the learning rate $l_{r}$, the input feature $x^{(l-1)}_{j}$, and the label $y_{\text{batch}}$ (see Eqs. 15 and 16). To reflect this, we adjust parameter trust using:
@@
    This formulation reflects the fact that both unreliable features and mislabeled data can strongly bias parameter updates. The use of min and max provides a conservative aggregation: trust cannot exceed the weakest evidence, and distrust must reflect the strongest warning. It ensures that trust in both input features and labels is explicitly propagated into the parameter trust.
@@
V-D Theoretical Properties of PaTAS
To ensure that PaTAS provides reliable and interpretable trust assessments, we first establish several fundamental theoretical properties describing its stability and logical consistency.
V-D1 Convergence of PaTAS
The convergence of the PaTAS is governed by the stability of its inputs and the structure of its recursive Parameter-Trust Update process. During training, PaTAS updates the internal trust opinions associated with network parameters using trust opinions on the inputs, labels, hyperparameters, and gradient information. This update mechanism is outlined in Alg. 1. As proved in Theorem 1, the PaTAS converges under specific situation. This convergence relies on some specific characteristic of the operators used to feedforward and revise the trust in the parameters $\theta$.
@@ V-D1 Convergence of PaTAS
Let $T_{\theta}^{(n)}$ denote the trust opinion at iteration $n$ for parameter $\theta$. If the revision of the trust in the weights is performed as specified in Alg. 1, then the PaTAS parameter will also converge.
@@
See Appendix C
V-D2 Symmetry and Invariance Properties of PaTAS
@@
Let $x=(b,d,u,a)$ be a binomial opinion with belief $b$, disbelief $d$, uncertainty $u$, and base rate $a$ where $b+d+u=1$. The opinion $\bar{x}=(d,b,u,a)$ is called the symmetric of $x$. Two binomial opinions $x$ and $\bar{x}$ are symmetric if they share the same uncertainty and base rate and have inverted belief and disbelief, i.e.,
$x=(b,d,u,a),\quad\bar{x}=(d,b,u.a)\quad\text{with }b+d+u=1.$
@@
These symmetry and invariance properties serve as fundamental consistency checks, ensuring predictable behavior under neutral, vacuous, or balanced evidence, while also simplifying evaluation by reducing the number of distinct trust scenarios that need to be considered.
@@ VI Evaluation and Results
The goal of our evaluation is to validate the PaTAS both theoretically and empirically. Specifically, we aim to demonstrate that PaTAS produces interpretable trust estimates that (i) converge during training under specific conditions, (ii) respect symmetry and invariance properties, and (iii) are able to provide interpretable assessments under realistic conditions such as noisy features, corrupted labels, or adversarial perturbations.
Our evaluation approach combines controlled synthetic degradations with real-world datasets. We systematically vary the trustworthiness of inputs ranging from fully trusted, fully uncertain, to fully distrusted, and observe how the created PaTAS propagates these trust assessments through the network. For each scenario, we track three complementary metrics: trust mass (belief), uncertainty mass, and distrust mass (disbelief). We also compare these values against standard model accuracy (after training) to understand how input trust affects output reliability.
We conduct three experiments of increasing complexity based on three different datasets:
@@ VI-A1 Experiment 1 - Breast Cancer Classification
For the evaluation, we degrade the training data in controlled ways and assign corresponding Subjective Logic trust assessments. We consider three extreme trust profiles (fully trusted $(1,0,0)$, fully distrusted $(0,1,0)$, and fully uncertain $(0,0,1)$) for both input features and label. These are combined (for inputs and label assessment) to form nine combinations. These dogmatic and vacuous opinions serve as canonical boundary cases in Subjective Logic: they express maximal trust, maximal distrust, and maximal uncertainty, respectively. We additionally include two intermediate scenarios, introduced below, to illustrate how PaTAS behaves under partial trust and partial distrust.
In practice, feature and label trust degradation may arise from poor-quality imaging, human annotation errors, or flaws in the preprocessing pipeline. In all experiments, we simulate such degradations using controlled perturbation functions. For fully uncertain feature opinion, we introduce additive uniform noise
$x^{\prime}=x+\mu,\quad\mu\sim U(-\eta,\eta),$
@@
$\eta=0.3\times\max(\text{features}).$
Each feature is perturbed with probability $0.3$. After noise addition, if $x^{\prime}$ lies outside the valid range $\min(\text{features}),\max(\text{features})]$, it is clipped to the corresponding boundary:
$x^{\prime}=\begin{cases}\min(\text{features}),&x^{\prime}<\min(\text{features}),\\ \max(\text{features}),&x^{\prime}>\max(\text{features}),\\ x^{\prime},&\text{otherwise}.\end{cases}$
To model distrust, we generate corrupted inputs by sampling from a uniform distribution over the feature space:
$x^{\prime}\sim U(\min(\text{features},\max(\text{features}),$
Label degradation is modeled analogously: uncertainty is introduced via random label noise, while full distrust is represented by a complete replacement of the labels.
Finally, to complement the boundary cases, we evaluate two intermediate trust scenarios:
@@
    the same scenario for a fully uncertain input features and fully uncertain labels, but where the assessment is set to $(0.25,0.25,0.5)$, reflecting partial distrust and trust, instead of a fully uncertain opinion (0,0,1).
@@ VI-A2 Experiment 2 - MNIST
Although these architectures are not state-of-the-art for MNIST, they are sufficient to evaluate the behavior of PaTAS across different model sizes and to demonstrate how trust propagates through the network. Each model uses the ReLU activation function for the hidden layer and Softmax for the output. Models are trained for 10 epochs with a batch size of 18 and a learning rate of 0.001, achieving test accuracies of 89%. We evaluate using fully uncertain Training data trust assessment functions for both input features and labels, corresponding to the case where we have no knowledge about the dataset.
@@ VI-A3 Experiment 3 - Poisoned MNIST
Implementation Details for All Experiments
During inference, multiplication operations in the feedforward phase are implemented using SL trust discounting, while addition operations employ a generalized SL averaging fusion operator to support summations over multiple inputs (see Definition 6). Trust revision uses the same averaging fusion. The PaTAS framework and Neural Network were implemented in Python 3.9 using NumPy. Two main modules were developed: PrimaryNN, which handles network structure, training, and inference, and PaTAS, which implements trust assessment and propagation functions defined in the operational flow (Fig. 7). These modules interact to dynamically evaluate and update trust during feedforward and backpropagation. The full implementation, including all modules, dependencies, and experiment scripts, is available as supplementary materials, with detailed instructions for reproducing all experiments.
@@ VI-B Results and Analysis
We focus on these three profiles because they represent the extreme and most informative boundary cases of input trust.
@@
All the results are depicted in Figures in Sections D-A, D-B, D-C, D-D, D-E, D-F, D-G and D-H (Appendix D) and summarized in Tables II, II, III and IV. They confirm several theoretical properties of PaTAS proven in Section V-D:
X Trust Assessment Y Trust Assessment $\epsilon=0.01$ $\epsilon=0.1$ Train (%) Test (%)
fully distrusted fully distrusted 0 0 63 38
fully distrusted fully uncertain 0 0 53 62
fully distrusted fully trusted 0 0 63 63
fully uncertain fully distrusted 0 0 96 3
fully uncertain fully uncertain 0.28 0.31 53 46
fully uncertain fully trusted 0.30 0.34 97 96.49
fully trusted fully distrusted 0 0 98 0.88
fully trusted fully uncertain 0.27 0.33 52 89
fully trusted fully trusted 0.79 0.90 98 99
(0.25, 0, 0.75) (0.25, 0, 0.75) - 0.38 71 90
(0.25, 0.25, 0.5) (0.25, 0.25, 0.5) - 0.17 53 46
@@
Hidden Neurons Trust $t$ Train (%) Test (%)
@@
For detailed results, we primarily focus on the three plots for each experiment) corresponding to feedforward of fully trusted input. Processing uncertain inputs naturally yields uncertain outputs, while symmetry implies that trusted and distrusted cases mirror each other. We also note that when processing fully trusted inputs, the distrust mass generally remains close to zero. Since trust, uncertainty, and distrust sum to 1, analyzing only the trust mass is sufficient in such cases. Therefore we keep only the trust mass after the training in Tables II, II, III and IV .
@@
We evaluated the system for two values of $\epsilon$ (0.1 and 0.01), as summarized in Table II. We recall that $\epsilon$ is the threshold used by the NodeTrust function in Alg. 1.
@@
In this experiment, PaTAS produces ten output trust opinions, one for each class. However they are almost the same so we record only one in the table. As shown in Table II, when features and labels are noisy, accuracy improves as model size increases. However, training accuracy remains consistently lower than test accuracy, with the gap widening for larger networks. This indicates that the models learn useful representations but cannot fully fit noisy training data, while still generalizing well on clean test samples. Test accuracy is consistently greater than training accuracy across all architectures, suggesting training noise makes the model underestimate its learning progress. Larger architectures yield more stable trust assessments, likely due to smaller average gradient magnitudes.
@@
As in the previous experiment, PaTAS has ten output trust opinions. These values are informative because the reliability of the inference path vary across labels, especially when some are poisoned and others are not. In Table III, we report the trust mass associated with label 3 (a clean class) and label 6 (a poisoned class; see Section VI-A3). Alongside overall train and test accuracy, we also report accuracy on clean test samples of digits 3 and 6, as well as accuracy on intentionally poisoned versions of those test samples.
The results suggest that PaTAS is able to reflect differences in reliability between the clean and poisoned classes, even under moderate corruption. Across patch sizes, the trust mass assigned to the clean label consistently remains higher than that assigned to the poisoned label, indicating a separation in how PaTAS evaluates the reliability of their respective inference paths. This separation depends strongly on the size of the patch used for poisoning:
@@
Patch size Trust for 3 Trust for 6 Train (%) Test (%) Clean 3 (%) Clean 6 (%) Poisoned 3 (%) Poisoned 6 (%)